OpenStudy (anonymous):

Grace is three years younger than her sister Jane. Twice the sum of their ages is 4 more than their father's age. if their father is 54, find the ages of Grace and Jane. (How do I start?)

OpenStudy (zzr0ck3r):

g = j -3 2(j+g)-4 = 54 can you solve this?

OpenStudy (anonymous):

it becomes j + g = 25

OpenStudy (anonymous):

Sorry I'm lost. Can you go step by step with me please?

OpenStudy (zzr0ck3r):

ok g = grace, j = jill grace is 3 years younger than her sister so if you take grace and add 3 to it you get her sisters age g+3 = j so g = j -3 ok?

OpenStudy (anonymous):

@zzr0ck3r how u will find the value of g or j

OpenStudy (anonymous):

as it becomes g = 25 - j or j = 25 - g

OpenStudy (anonymous):

Okay I get it. Thanks everyone!

OpenStudy (zzr0ck3r):

twice the sum of there ages is 4 more than ther fathers so twice there ages added together equals there fathers age + 4 2(g+j) = 54+4 so we have g - j = -3 and 2g+2j = 58 or g+j = 24 g-j = -3 g+j = 29 add the two equations 2g = 26 g = 13 13+j = 29 j = 16

OpenStudy (zzr0ck3r):

make sense @Ekaansh ?

OpenStudy (anonymous):

k i forgot about the second eq. now we can solve it by any method :)

OpenStudy (zzr0ck3r):

word

OpenStudy (anonymous):

[(j-3) + g] 2 - 4 = 54 ?

OpenStudy (zzr0ck3r):

what is your queston @li_za12 ?

OpenStudy (anonymous):

is that right?

OpenStudy (zzr0ck3r):

you cant solve 2 variables in one equation. but yes this would work for one equation, its the same thing I did.

OpenStudy (jiteshmeghwal9):

Let the age of grace be \(x\).then his sister's age=\(x+3\) A/q, \(2(x+(x+3))=54+4\) \(2(x+(x+3))=58\) solve this & gt ur answers :)